the sum of the page numbers on consecutive pages of a book is 267. what are the page numbers?: the sum of the page numbers on consecutive pages of a book is 267. what are the page numbers?
Answer by checkley77(3624)

x+(x+1)=267
2x=267-1
2x=266
x=266/2
x=133 for the first page number & 133+1=134 for the second page number.
Proof:
133+134=267
267=267